### v1

# Confining quantum field theories

Abstract
It is widely believed, and axiomatically postulated in mathematical quantum field theory, that the vacuum is a unique vector state (up to a phase factor). The recent solution of the quantum Yang-Mills theory of the strong interaction revealed the presence of two vacua and a mixed quantum state. The second, confining vacuum, is an eigenstate of an auxiliary field, with a non-zero eigenvalue, as opposed to the zero eigenstate of the perturbative vacuum.
Here, I show that this non-trivial vacuum structure is characteristic of confining quantum field theories, and implies confinement, in the sense that vacuum expectation values between states separated at large, space-like distances, tend to zero, whereas in ordinary quantum theories with a unique vacuum, they are known to satisfy the cluster decomposition principle, and at large separations tend to free, asymptotic states.
I conclude with a discussion and some tentative comments regarding the general status, problems, and interpretation of quantum field theory for the four known interactions.
1 Introduction
In [1], I used a specific set of Feynman rules, in order to satisfy Gauss’s law and estimate the interaction energy between two static sources in Yang-Mills theory. After using an auxiliary operator, a Lagrange multiplier field, $\lambda$, the effective action derived, although at a particular Lorentz frame, was shown in [2] to admit two distinct, stable vacua, the perturbative vacuum, $|\Omega_{0}>$, with $\lambda=0$, and the confining vacuum, $|\Omega_{\mu}>$, with $\lambda^{2}=\mu^{2}$, with a dimensionful constant, $\mu$, generated via the Coleman-Weinberg mechanism. Physics at each vacuum, as well as in the mixed vacuum state, is Lorentz-invariant. In [3] the role of the confining vacuum as an eigenstate of the gauge-invariant auxiliary operator $\lambda^{2}\,|\Omega_{\mu}>\,=\mu^{2}\,|\Omega_{\mu}>$, was explained, and contrasted with the ordinary, conventional quantum field theory, with the unique vacuum postulate, which is equivalent to the cluster decomposition property. A related feature of the confining theory is the fact that $\lambda^{2}$ is a non-trivial operator, that commutes with all other operators of the theory, but is not identically zero, or a multiple of the identity.
Here, I show that this vacuum structure of the quantum Yang-Mills theory is characteristic of all confining theories, and implies confinement, in the sense that all correlation functions at large spacelike separations tend to zero, instead of satisfying the cluster decomposition property (another unwarranted postulate of conventional quantum field theory, equivalent to the existence of a unique vacuum).
Also, in the case of Yang-Mills theory, there is a vacuum energy density difference between the two vacua that depends on $\mu^{2}$, the non-zero eigenvalue of the non-trivial operator, $\lambda^{2}$. I explain how the approach to the zero limit of the aforementioned correlations is related to this energy, but I also note that, in the general case, the non-trivial operators and vacua do not have to correspond to an energy difference. Confinement is a general characteristic of the non-trivial structure with multiple vacua, and all states in such theories are confined.
In theories where there is such an energy difference between different vacua, however, as explained in [3], there is no Lagrangian description of the theory; if such a gauge- and Lorentz-invariant description existed, one would be able to connect the two vacua using operators of the Lagrangian, and thus lower the energy, which is not possible, by the definition of a vacuum. In general confining theories, with a mixed vacuum state, the absence of asymptotic states also implies the absence of an S-matrix, and makes traditional interpretations [4, 5, 6] even more difficult.
In Sec. 2, I review the solution of Yang-Mills theory. In Sec. 3, I show the general confining property of such theories, and in Sec. 4, I make some general, more or less tentative comments, regarding the general status of quantum field theory and the other physical interactions.
2 The Yang-Mills theory
The effective action that incorporates Gauss’s law in Yang-Mils theory, with an auxiliary Lagrange multiplier field, $\lambda$, was derived as
$S_{\rm eff}=\int\frac{1}{2}{E_{i}^{a}}E_{i}^{a}-\frac{1}{2}{B_{i}^{a}}B_{i}^{a}+\lambda^{a}\,{D_{i}}{E_{i}^{a}}+U(\lambda),$ (1)
where $E^{a}_{i}=F^{a}_{0i},B^{a}_{i}=-\frac{1}{2}\epsilon^{ijk}F^{a}_{jk}$, $F^{a}_{\mu\nu}=\partial_{\mu}A^{a}_{\nu}-\partial_{\nu}A^{a}_{\mu}-g_{c}f^{abc}A^{b}_{\mu}A^{c}_{\nu}$, for the non-Abelian gauge group with generators $T^{a}$, structure constants $f^{abc}$, and coupling $g_{c}$, written in terms of the fields $A_{\mu}=T^{a}A^{a}_{\mu}$ , $\lambda=\lambda^{a}T^{a}$.
The effective action is gauge invariant under a local gauge transformation $\omega(\alpha)=e^{iT^{a}\alpha^{a}(x)}$, $A_{\mu}\rightarrow\omega A_{\mu}\omega^{-1}+\frac{i}{g_{c}}\omega\partial_{\mu}\omega^{-1}$, $\lambda\rightarrow\omega\lambda\omega^{-1}$.
The effective potential $U(\lambda)$ is of the Coleman-Weinberg form, with a minimum at $\lambda^{2}=\mu^{2}$, at a generated scale, $\mu$, but appears inverted in the effective action ($\lambda^{2}=\lambda^{a}\lambda^{a}$).
Thus, the inverted potential, $-U$, has a local minimum at $\lambda=0$, and a global maximum at $\lambda^{2}=\mu^{2}$, but because of Gauss’s law and the presence of gauge kinetic terms, the analysis of [2] demonstrated that they are both stable vacua, and in [3] it was shown that they should be interpreted as eigenstates of the auxiliary, gauge-invariant operator, $\lambda^{2}$.
This is a new mechanism of scale generation. Although the combinatorics of the Coleman-Weinberg effective potential are well known, arising from the infrared singularities of the massless, self-interacting, gauge fields, here there is no spontaneous symmetry breaking, $\lambda$ is an auxiliary field, without a kinetic term, and without any additional degrees of freedom, and $\mu^{2}$ is an eigenvalue, not a vacuum expectation value from spontaneous symmetry breaking.
$|\Omega_{0}>$, with eigenvalue $\lambda=0$, is the perturbative vacuum, with the usual Coulomb interaction, and $|\Omega_{\mu}>$, with $\lambda^{2}|\Omega_{\mu}>=\mu^{2}|\Omega_{\mu}>$, is the confining vacuum.
A rich structure of the proposed Feynman rules and effective action was already derived at previous works, including the confining interaction, soliton solutions connecting the two vacua, as well as the possible solution of the strong-CP problem, a first-order phase transition at high temperatures, and chiral symmetry breaking after the inclusion of fermions. These will be further investigated in future works. Here, as described before, I will focus more on the vacuum structure of Yang-Mills theory, and its possible generalizations.
First, it should be noted that there is an energy density difference between the two vacua: the vacuum energy density of $\Omega_{0}$ is zero, and the vacuum energy density of $\Omega_{\mu}$ is positive (equal to $-U(\mu^{2})$). The canonical formalism in [2] derives the energy (Hamiltonian) as
$H=\int d^{3}x\left(\frac{1}{2}E_{i}^{a}E_{i}^{a}+\frac{1}{2}B_{i}^{a}B_{i}^{a}-U\right).$ (2)
However, since both vacua are stable, there is no well-defined Lagrangian or energy-momentum tensor that connects the two vacua. If there was, it would be possible to perturb one vacuum, using operator configurations from the Lagrangian, and lower the energy of one vacuum, which is not possible. The effective action of (1) is given at a particular Lorentz frame, in order to express the Gauss’s law constraint and derive the force between two static sources. The physics at each vacuum, however, as well as overall, is Lorentz invariant. At the vacua there is a well-defined energy-momentum tensor, $T^{\mu\nu}=T^{\mu\nu}_{\rm YM}-g^{\mu\nu}U$, where $T^{\mu\nu}_{\rm YM}$ is the usual energy-momentum tensor for perturbative Yang-Mills, and $g^{\mu\nu}$ is the Lorentz metric.
The canonical analysis also showed that the auxiliary operator, $\lambda^{2}$, is also a non-trivial operator that commutes with all other operators of the theory, without being identically zero, or a multiple of the identity. The associated non-trivial, non-unique vacuum structure, and the decomposition into different, distinct vacua, that are eigenstates of this operator, was shown to correspond to fundamental constructions of operator algebras, namely the Gelfand-Naimark-Segal (GNS) theorem.
Also, the breakdown of the cluster decomposition principle, which is equivalent to the uniqueness of the vacuum, was shown to be related to the above structure, as well as the confining properties of the theory. In the next Section, I show that confinement is characteristic of every theory with a similar, non-trivial, vacuum structure, with different vacua that are eigenstates of a non-trivial operator that commutes with every other operator of the theory.
3 Vacuum state and confinement
Quantum field theory deals with local operators, $Q$, defined by smeared out fields, $\Phi$,
$Q=\int f(x)\,\Phi(x)\,d^{4}x,$ (3)
with smearing functions like $f(x)$ above, that have support (take non-zero values) at finite regions of four-dimensional Minkowski space-time, for example, of the form $V\times T$, that describe physical operations, measurements of observables, that are performed in the system, at the finite three-dimensional volume, $V$, over a period of time, $T$.
Pure states of the physical system are described by normalized vectors, $|\Psi>$, of a Hilbert space, and probabilities for physical observations at such pure states are given by inner products, $<\Psi|Q|\Psi>$, whereas mixed states are described by density matrices.
Translated operators are given by
$Q(x)=U(x)\,Q\,U^{-1}(x),$ (4)
with the unitary $U(x)=e^{iP_{\mu}\cdot x^{\mu}}$, where the operator $P_{\mu}$ is constructed from the $T_{\mu\nu}$ at each vacuum.
One can then define operators like [5]
$\tilde{Q}(p)=\int Q(x)e^{-ip\cdot x}d^{4}x,$ (5)
which changes the energy-momentum of a state vector by $p$, and
$Q(f)=\int Q(x)f(x)d^{4}x,$ (6)
which shifts the energy-momentum of a state vector by the support, $\Delta$, of $\tilde{f}(p)=\int e^{ip\cdot x}f(x)d^{4}x$. Then, if the spectral supports of the state vectors $|\Psi_{1}>$ and $|\Psi_{2}>$ are the regions $E_{1}$ and $E_{2}$, respectively (also subsets of the four-dimensional momentum space), we have
$<\Psi_{2}|Q(f)|\Psi_{1}>=0\,\,\,\,\,{\rm if}\,\,\,\,(E_{1}+\Delta)\bigcap E_{2}=\emptyset.$ (7)
The vacuum of the theory, a vacuum state, $|\Omega>$, is defined by the property that $<\Omega|Q^{*}(f)Q(f)|\Omega>=0$, for any operator, $Q(f)$, that lowers the energy (the support of $\tilde{f}(p)$ is not in the positive cone, $p^{0}>0$, $p^{2}>0$).
It is usually assumed that the vacuum is a unique (up to a phase) vector state. However, the recent solution of Yang-Mills theory, demonstrated the existence of the two aforementioned vacua, that are eigenstates of the operator $\lambda^{2}$, which commutes with all other operators of the theory.
I will show here that this structure implies confinement in a rather general way, such that, for any two local operators, and any configuration of spacelike separated points at equal time, $x_{1}=(t,\vec{x}_{1})$, $x_{2}=(t,\vec{x}_{2})$,
[…]
at large spacelike separation $R=|\vec{x_{1}}-\vec{x_{2}}|$.
This is opposed to the situation of ordinary, conventional, quantum field theory, with a unique vacuum vector, $|\Omega>$, where the limit in the right-hand side of (8) is a product of free states
$<\Omega|Q_{1}(x_{1})Q_{2}(x_{2})|\Omega>\,\,\,\rightarrow\,\,<\Omega|Q_{1}(x_{1})|\Omega><\Omega|Q_{2}(x_{2})|\Omega>.$ (9)
Obviously, (9) is the cluster decomposition property, that expresses the fact that two interacting states of a system, when separated at large distances become non-interacting. This is true for theories like quantum electrodynamics, but cannot possibly be expected to hold for confining theories, like the strong interaction.
First, one notes that any operator, $Q$, can be decomposed in three parts, as $Q=Q^{-}+Q^{+}+Q^{0}$, that satisfy $Q^{-}\Omega_{\mu}=0$, $Q^{+*}\Omega_{\mu}=0$, and with the spectral support of $Q^{0}\Omega_{\mu}$ being a bounded subset of the forward and backward light cones of the energy-momentum space with $|p^{0}|\leq c$, for any arbitrary constant, $c$.
Indeed, after defining $\tilde{Q}(p)=\int f(x-y)\Phi(x)e^{-ip\cdot y}d^{4}y\,d^{4}x$, for any $Q=\int f(x)\Phi(x)d^{4}x$, and using a partition of unity, $F_{+}+F_{-}+F_{0}=1$, with
$F_{+}(p)=0\,\,{\rm for}\,\,p^{0}<b,$
$F_{-}(p)=0\,\,{\rm for}\,\,p^{0}>-b$
$F_{0}(p)=0\,\,{\rm for}\,\,|p^{0}|>c$
$0<b<c,$ (10)
$Q=\int(F_{+}(p)+F_{-}(p)+F_{0}(p))\tilde{Q}(p)d^{4}p$ satisfies the criteria for the required splitting.
Then
$<\Omega_{\mu}|Q_{1}(x_{1})Q_{2}(x_{2})|\Omega_{\mu}>=$
$<\Omega_{\mu}|(Q^{-}_{1}(x_{1})+Q^{0}_{1}(x_{1}))(Q^{+}_{2}(x_{2})+Q^{0}_{2}(x_{2}))|\Omega_{\mu}>,$ (11)
and for large $R$ one can commute $Q^{+}_{2}$ to the left and $Q^{-}_{1}$ to the right, where they give zero. The approach to zero from the commutators will be discussed later.
The remaining term is $<\Omega_{\mu}|Q^{0}_{1}(x_{1})Q^{0}_{2}(x_{2})|\Omega_{\mu}>$, and the spectral supports, $E_{1}$, $E_{2}$, of the vectors $|\Psi_{1}>=Q^{0}_{1}|\Omega_{\mu}>$, $|\Psi_{2}>=Q^{0}_{2}|\Omega_{\mu}>$ are as described above, bounded subsets of the energy-momentum light-cones.
Since $|\Omega_{\mu}>$ is an eigenvector of the operator $\lambda^{2}$, with a non-zero eigenvalue, and this operator commutes with any $Q$, this remaining term is also proportional to $<\Omega_{\mu}|Q^{0}_{1}(x_{1})\lambda^{2}(f)Q^{0}_{2}(x_{2})|\Omega_{\mu}>$.
Then, for spectral supports, $E_{1,2}$, that are bounded subsets of the light cones, it is easy to see that one can pick a function, $f$, with spectral support, $\Delta$, such that $(E_{1}+\Delta)\bigcap E_{2}=\emptyset$, and use (7), thus giving the final zero result in (8).
It is also interesting to notice the rough resemblance of this proof to the Feynman diagram of Fig. 5 of [2], which was used in order to derive the confining interaction.
The neglected terms from the commutators at large spacelike distances decrease to zero, as in ordinary field theory, at least at the order of $1/R$. In the case of Yang-Mills theory, with the confining vacuum having an energy density of order $\mu^{4}$, and for states localized in a physical volume of order $V$, terms of order $e^{-mR}/R$ are expected, with $m\sim\mu^{4}\,V$. For pure Yang-Mills, however, without fermion fields or other scales, $V$ is arbitrary, and a Coulomb (or Luscher-type) term cannot be excluded. Strictly speaking, in the sense of conventional field theories, there is no “mass gap”, at least not in this proof ($c$ is arbitrary) but there is an energy density difference for all states on the confining vacuum, as described before.
The important point of this work, however, is the result (8), which is a general confining statement, for any theory that has a similar structure of non-trivial vacua, and commuting non-trivial associated operators $\lambda$ (there may be more than one) with vacuum eigenstates. It is to be contrasted with the standard, unwarranted assumption (9) of conventional quantum field theories, that cannot possibly hold in confining situations.
In other confining theories, these related operators $\lambda$ do not all have to be related to an energy density difference. In some theories, therefore, unlike the strong interaction, there may exist a Lagrangian formalism and still have confinement; these and other problems of interpretation are further discussed in the next Section.
Finally, it is obvious that a similar argument holds for any number of spacelike separated operators, $Q$, as in ordinary quantum field theory.
4 Quantum field theories, old and new
Conventional quantum field theories [4, 5, 6] (typical examples being quantum electrodynamics with a coupling $e^{2}$, or a single scalar field, $\Phi$, with quartic coupling $g\Phi^{4}$ and without symmetry breaking) consist of a unique vacuum state, $|\Omega>$, and operators, $Q$, localized in regions of spacetime. Physical states are represented by vectors, $|\Psi>$, of the Hilbert space, $H$, of the theory, and mixed states by density matrices. Every vector state, $|\Psi>$, of the physical system of such theories can be built (approximated) by operators acting on the vacuum (the vacuum is cyclic). That is, one can find operators, $Q$, such that $|\Psi>\,\approx Q|\Omega>$, at any level of approximation, for any $|\Psi>$ (the set of all such $Q|\Omega>$ is dense in $H$).
It is easy to see then, that in such theories, any operator $\lambda$, that commutes with all other operators of the theory, is trivial (either zero or a multiple of the identity).
The structure of the vacuum, even in these theories, however, is far from trivial (the main reason being the cyclicity property). Physically, the vacuum state contains all the vacuum-to-vacuum diagrams and transitions, all closed loop Feynman diagrams, so it is different for different values of the coupling, and accordingly, the Hilbert space of the theory depends on the coupling and is different from the “free” theory vacuum with zero coupling (Haag’s theorem). Free, asymptotic states, however, can be built, and an $S$-matrix can be defined, essentially based on the cluster decomposition property, which is equivalent to the uniqueness of the vacuum.
The relation of these two properties, which is, actually, a mathematical, logical equivalence (uniqueness of the vacuum $\Leftrightarrow$ cluster decomposition property) was observed as early as in [7], it is noted in [4], and is also mentioned in [5, 6] in connection with results from operator algebras, such as the GNS construction. All these, as well as subsequent works, however, and eventually quantum field theory textbooks, adopt them as a postulate (a mathematical axiom) in order to build and justify the $S$-matrix construction, which was used to describe the experimental results investigating particle physics. In fact, the early related theorems [7] were derived even before the proposal of the quark model, which was later included in the Standard Model, and its eventual establishment. The weak interactions also, after the spontaneous symmetry breaking mechanism, are expected to have a unique vacuum state, their excitations become massive with vacuum expectation values of Higgs fields, and the related interactions diminish as Yukawa terms, also giving the possibility of an $S$-matrix construction.
The strong interaction, however, with a linearly rising interaction, and confinement, cannot possibly be expected to satisfy the cluster decomposition property, and thus lead to a similar construction. There are no asymptotic states, and there is no such thing as an $S$-matrix. Accordingly, the vacuum state is not unique, there is a non-trivial operator, such that the second (confining) vacuum is an eigenstate of this operator, leading to confinement properties like (8), instead of (9). This is a consequence of the non-linear, self-interaction of the massless, spin-1, gauge bosons.
Although many features of the strong interaction can be seen and investigated with the methods presented here and in previous works, it is stressed that many properties of ordinary quantum field theory do not hold in such theories. Along with the breakdown of the cluster decomposition principle, one has the absence of an $S$-matrix formalism, and even the absence of a globally defined Lagrangian description. Physics is described locally, in terms of local operators in different compact regions of space-time, and can be matched and patched together, for example, in the spirit of a Wilsonian renormalization group, but globally, one does not have the luxury of an $S$-matrix or Lagrangian.
Once we dismiss the Lagrangian formalism as fundamental, we can investigate both new and conventional quantum field theories also in a new light (and also justify and revisit some older interpretations). The physical quantities that are measured are the probabilities that two or more states in a given volume interact; this defines a coupling constant together with a corresponding interaction operator, again in the spirit of the renormalization group, and the conditions under which these operators are patched together in a global Lagrangian can be investigated; singularities, blow-ups, and Landau poles, then may not be considered as problems of the theory per se, but of the Lagrangian construction (much like the infinities of early quantum field theory were remedied with the process of renormalization). Starting by postulating a Lagrangian may be useful and impressive in expressing, predicting, and justifying symmetries of the theory, but is like starting from the conclusion; the problem is how, if, and under what limits and conditions, this formalism can be reached.
On a mathematical sidenote, observables are elements of operator algebras, $A$, states are specifically defined (normalized, etc) elements of the dual space, $A^{*}$, and interactions are (also appropriately and consistently defined) elements of the double dual, $A^{**}$. The natural embedding of a space into its double dual, gives the possibility of defining interactions also as operators (and even there may be leeway; these spaces are not always reflexive). The various “coupling constants” then, are merely normalizations and initial values of elements of $A^{**}$ and, of course, are not constant, as is already well-known in quantum physics. These mathematical operations of taking the dual spaces, and the associated weak topologies involved, may seem “high-brow” or esoteric, but are, actually, the way that the physical measurements and operations are performed [6].
The weak interaction is also non-Abelian, but after spontaneous symmetry breaking, the Coleman-Weinberg potential that is generated by the mechanism of [1] does not admit a second, non-zero extremum, and does not change the unique vacuum structure, although a full treatment is also necessary.
It is also well-known, that the $S$-matrix in these conventional theories with a unique vacuum, usually starts with a tree contribution that is essentially a classical scattering result, and the role of relativistic, quantum physics is mainly to calculate higher order corrections in this result. In new, non-trivial, confining theories with multiple vacua, the phenomena are intrinsically quantum, and new experimental, as well as theoretical methods need to be devised.
It should also be stressed that, although the quantum electromagnetic and weak interactions are expected to admit a unique vacuum and a well-defined $S$-matrix, any other theory that includes the strong interaction, unified or not, including the Standard Model and its extensions, does not admit a conventional interpretation, except maybe in some approximations, that also need to be specified. An interesting discussion of relative scales and confining theories at various limits appears in [8], although in a different context.
It seems very likely, that the fourth physical interaction, gravity, also highly non-linear, would admit a similar, non-trivial structure, with multiple and mixed vacuum states, and confinement properties related to cosmological and black hole horizons, inflation, information puzzles, even a quantum version of Mach’s principle. One can start in a similar manner and consider the constraints of the gravitational interaction, utilize auxiliary fields in order to express them and investigate the generation and stability of new vacua. A more daring possibility, that is perhaps worth some investigation, would start with the realization that the absence of a conventional quantum field theory formalism and an $S$-matrix, also nullifies the traditional theorems that utilize its properties and symmetries (some versions of them may hold, of course, but need to be re-examined). One may then attempt to relate, couple, or somehow connect, an intrinsic symmetry (like the non-Abelian, weak interaction, or a new such interaction) to a subgroup of the Lorentz symmetry group of the spin-2 field. Instead of a Higgs mechanism of mass generation via spontaneous symmetry breaking with a vacuum expectation value, then one would have a quantum generation of a mass scale as an eigenvalue of a new vacuum state and, as a bonus, include gravity in the process.
As is usually stated in a conclusion, these physical and mathematical investigations will be considered in future works.